Prev Next

SW&IT Meeting Minutes 2019-11-04

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191021
LR20191021 - Bec, Matthieu 61.0
LR20191021 - Holmberg, Carl 40.0
LR20191021 - Colson, Andrew 40.0
LR20191021 - Edmeades, David 40.0
LR20191021 - Sola, Igor 40.0
LR20191021 - Kubanek, Petr 76.5
LR20191021 - Hooper, Stephen 80.0
LR20191021 - Zhang, Yang 64.0
LR20191021 - Zhang, Xianyu 80.0

News from management

  • Steward building construction Dec 4 - Jan 12
  • Tucson CR construction Nov 4 - 5
  • Staff announcement
  • 10/11 is a holiday, file next week reports on Friday.

Previous Action Items

Open topics

  • SPIE 2019 conference?
    • Nov 13 deadline submit abstract
  • TNO visit Nov 13 to showcase their next-gen AdSecs
  • VPN moutain + ALTA ?
  • 4G access to mountain, lack of clear guidelines, instructions (where is the hotspot)

  • Valley telecom contact - use MGIO on-call number (SH will update callout list maintained by Louise)

Team reports

MB

  • OC, WPC, Management
  • finishing up dms repo browser
  • UITS coloc meeting and egress
  • sciarc meeting
  • mostly sick
next
  • xxx

PK

  • wiki improvements (tcs start)
  • merged 2019B pull requests, preparing for final testing
  • patch for IT7903 - shows proxy name
  • WPC meeting
  • git difftool -x vimdiff -y
  • SSD lessons meeting
  • telescope power-up support
  • IT7909 - LSS lost connection to RPCserver
  • AZCam C++11 class (for WFS, GuideCam)

XZ

Last week:
  • 4D SW support
  • SOUL VM and PC dicussion
  • check analyze VM
  • start AO-data document
  • data copying for jumping actuator analyze demonstration

This week:
  • Join SOUL commission night observation (Wednesday-Saturday)
  • analyze VM
  • conitune AO-data document
  • jumping actuator analyze demonstration

DE

  • Final decision on colo? Double-ended NAT really not an option?
  • Server room cleanup for alarm work
  • Peter Labview provisional success
  • Converging on appropriate model of storage to present; should have numbers this week
  • VT outage
    • Do they have a way to push alerts to us?
    • Science requests to be on the list for the UITS Solarwinds alert on the MGIO router
    • Wireless NIC in one or more obs machine to bridge hotspot and LBTO network
  • VM hardware shuffling

CH

Last week:
  • DMS:

    • Updated imonnit data description file to include recently added sensors.
    • Reviewed several self-hosted URL shorteners, began testing a Node.js-based example.
  • SupportScripts: RB_Science.py bug fixed, was opening new FITS files before external copy process was complete.

This week:
  • ALH: Integrate ALS script/config file changes.
  • SupportScripts: Follow up with John Nill on changes.

YZ

  • Fixed a bug in the LBC/MODS ingestion script for web-dms.
  • Working on patching the TCS IIF C interface to address the LBC non-sidereal guiding problem (a first step).
  • Work with Olga to devise some tests for MODS GUI problem (IT#7855, tentatively on Monday).
  • Working on redesigning the AO-Cravat database as well as the web interface (need to discuss with Al Conrad).
  • For AO-Cravat, Strehl calculation is still buggy, and FWHM calculation is missing. Need to discuss with Steve Ertel about them.
  • New safety training classes.

AC

  • argos-dx navigation menu organization - TODO: review with operators  
  • argos jenkins build - testing against different builds of TaN
next  
  • create systemd for pms-lasair  
  • add misc scripts to github - display scripts (facsum.sh, etc)
  • OAC - automate commit of oacontrol.conf

IS

Last week:
  • miniQ
    • Implemented first version.
This week
  • miniQ
    • Hope to field a version of it tonight and work on improvements, particularly the report part.

SH

  • VM1 disk internal disk array built
  • ao-analysis on VM1
  • Colocation meeting with UITS
  • Gustavo and argos remote desktops
  • Move VMs from VM2 to VM1 mountain
  • Power-failure stuff
  • network badness
  • mountain archive badness
next:
  • Betsy and google group changing email.arizona.edu
  • mountain archive -- what is going on, and what can I do?
  • old vm1 rebuild
  • Tuesday up the mountain
    • wire up new VM1 better
    • mountain emergency response device
  • old vm2 out of vCenter, new vm1 into vCenter
    • vCenter upgrade
    • VMWare licensing?
Topic revision: r15 - 04 Nov 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback